What is 100010002 - 00011011002

1
Expert's answer
2021-06-29T09:37:28-0400
"\begin{matrix}
 & \ \ \ 10001000_2\\
 - \\
& 0001101100_2\\
\hline
& \ \ \ \ \ \ 0011100_2\\

\end{matrix}"

Or


"10001000_2=8+128=136"

"0001101100_2=4+8+32+64=108"

"136-108=28"

"28=11100_2"